OpenStudy (rayford):

Which of the following is a factor of 1 + 27x^3? 1 + 3x 1 + 3x + 9x2 1 + 6x + 9x2 None of the above

OpenStudy (neo92):

\[1+27x^{3}\] =\[1^{3}+27x{3}\] =\[(1+3x)(1-3x+9x{2})\]

OpenStudy (neo92):

therefore the answer for your question is (1+3x)

OpenStudy (rayford):

thank you i understand now!!

OpenStudy (neo92):

u r welcome.. by the way in genaral a^3 + b^3 = (a+b)(a^2 - ab + b^2) a^3 - b^3 = (a-b)(a^2 + ab + b^2)

OpenStudy (rayford):

sir yes sir!